# Simple Semi-supervised Knowledge Distillation from Vision-Language Models via Dual-Head Optimization

### v1

@@ 4.2 Experimental Results
[F1]: Effectiveness of DHO compared to SHO.
We first compare DHO with SHO on ImageNet under 1, 2, 4, 8, and 16-shot settings using ResNet-18 (trained from scratch) and self-supervised ResNet-50. As shown in Table 1, DHO consistently outperforms SHO, with further gains from DHO-F that incorporates few-shot teachers. We also compare against dual-head KD methods on ResNet-50: SSKD and DHKD achieve 55.2/58.1/60.0/62.3/64.0 and 25.6/34.8/42.7/49.2/55.2, respectively, while DHO surpasses both across all shot counts. To assess generality, we evaluate on 10 additional datasets using ResNet-18. DHO consistently outperforms SHO, with larger improvements from few-shot teachers (Figures 5 and 6). Results for MobileNetV2 are shown in Figures 12 and 12. Table 2 shows these gains incur minimal overhead in FLOPs and throughput, especially on ImageNet and negligible for smaller datasets with fewer classes ($C<1000$). See inference overhead for MobileNetV2, ViT-B/16, and ViT-L/14 in Table 7.
@@
CE+KD (feature) [101] 62.3 85.0
CE+KD (logit) [13] 66.2 88.8
@@
[F2]: Enhanced feature representation of DHO.
To validate our claim that mitigating gradient conflicts improves feature representations, we evaluate features using the standard linear evaluation protocol [13]. We train CE+KD (feature), CE+KD (logit), and DHO under the 16-shot semi-supervised setting on ImageNet, freeze the feature extractor $g$, and train a new prediction linear head $h_{\text{LE}}$ on top of $g$ using fully labeled data. As shown at the top of Section 4.2, DHO achieves higher Top-1 and Top-5 accuracy than other methods. To further assess feature quality, we visualize embeddings $z$ using t-SNE [89] in Figure 7. Compared to the CE+KD (logit) baseline, DHO produces more compact and class-separated feature clusters. These results support our claim that DHO enhances feature representations by mitigating gradient conflicts; this improvement leads to better performance compared to SHO, as discussed in [F1] and illustrated in Table 1, Figures 5, 6, 12 and 12.
Figure 8: Results of ablation studies on dual-heads interpolation strategy in Eq. 11 of DHO.
Figure 9: Qualitative results on challenging cases.
[F3]: Effectiveness of dual-head interpolation.
We evaluate the effectiveness of dual-head interpolation (Eq. 11) by comparing DHO with CE+KD (logit), CE+KD (feature), and ablations DHO ($h_{\text{CE}}$) and DHO ($h_{\text{KD}}$), which predict using only one head, i.e., $\alpha=1$ or 00. We use ResNet-50 for ImageNet and ResNet-18 for the other datasets. As shown in Figure 8, DHO outperforms DHO ($h_{\text{CE}}$) by an average of 1.6% across 11 datasets, with a maximum gain of +3.4% on ImageNet and no degradation on any dataset. Since $\alpha$ and $\beta$ are inference-time hyperparameters, dual-head interpolation introduces minimal overhead while consistently improving or maintaining performance. We further conduct case studies on its effect. Figure 9 illustrates three challenging examples: CE head ($h_{\text{CE}}$) is correct in the first, KD head ($h_{\text{KD}}$) in the second, and both fail in the third case, yet the proposed combined prediction is correct—demonstrating the ability of DHO to resolve individual head failures. See additional analysis in Section E.3.
Method Architecture Params 1% 10%
(M) (%) (%)
Self and Semi-supervised Learning
MSN [3] ViT-B/4 86 75.7 80.2
Semi-ViT [7] ViT-L/14 307 77.3 83.3
Semi-ViT [7] ViT-H/14 632 80.0 84.3
CLIP-based Training
CLIP [51] ViT-B/16 86 74.3 80.4
REACT [55] ViT-B/16 86 76.1 80.8
REACT (Gated-Image) [55] ViT-B/16 129 77.4 81.8
CLIP [51] ViT-L/14 304 80.5 84.7
REACT [55] ViT-L/14 304 81.6 85.1
REACT (Gated-Image) [55] ViT-L/14 380 81.6 85.0
Co-training based Methods
CT [74] Multi-arch (2) 608 80.1 85.1
MCT [74] Multi-arch (2) 608 80.7 85.2
CT [74] Multi-arch (4) 1071 80.0 84.8
MCT [74] Multi-arch (4) 1071 80.5 85.8
Knowledge Distillation
SimCLR v2 distill [13] ResNet-50 (2×+SK) 140 75.9 80.2
SimCLR v2 self-distill [13] ResNet-154 (3×+SK) 795 76.6 80.9
DHO (ours) ViT-B/16 86 81.6 82.8
DHO (ours) ViT-L/14 304 84.6 85.9
Zero-shot VLMs
SigLIP [112] ViT-SO400M/14 400 83.1 -
DFN [22] ViT-H/14 632 84.4 -
Table 5: Results of ablation studies on language-aware initialization and KD-head alignment for VLM students on ImageNet with 1% labeled data.
Init ($h_{\text{CE}}$/$h_{\text{KD}}$). Align. Accuracy (%)
@@
Figure 10: Visualization of grid search results for $\alpha$ and $\beta$ using ResNet-50 on ImageNet under the 16-shot semi-supervised setting.
[F4]: DHO achieves SoTA performance on ImageNet under low-shot settings.
We compare DHO to previous state-of-the-art (SoTA) methods on ImageNet under 1% and 10% labeled data. All results are taken from published papers, except for DHO. As shown in Section 4.2, DHO with ViT-L/14 surpasses the previous SoTA by 3% (1% data) and 0.1% (10% data), while using 218M and 767M fewer parameters, respectively. Notably, DHO achieves 81.6% accuracy with only 86M parameters—the same as REACT [55] using 304M. These results show that DHO is a simple yet effective method.
Effectiveness of language-aware initialization and KD head alignment.
Table 5 shows ablation results on language-aware initialization (Init.) and KD-head alignment (Align.) on ImageNet (1%). We use ViT-B/16 student and ViT-L/14 teacher for computational efficiency. We observe that applying Init. on each head independently improves performance, and applying Align. further enhances it.
@@
To understand how $\alpha$ and $\beta$ in DHO affect performance, we visualize grid search results on ImageNet with ResNet-50 under the 16-shot setting. Figure 10 shows performance peaks at balanced heads ($\alpha\approx 0.5$) and degrades at extreme values ($\alpha\approx 0$ or $1$) regardless of $\beta$. With balanced heads, performance remains stable for $\beta\in[0.1,1]$, optimizing around $\beta=0.3$. Furthermore, under mild assumptions, Appendix A shows that DHO $\epsilon$-approximates SHO w.r.t the $\ell_{1}$ norm by setting $\alpha=\lambda$ and $\beta=\tau$, where $\tau$ is the corresponding temperature hyperparameter for $\beta$ in SHO. Notably, $\lambda$ and $\tau$ are training hyperparameters, while $\alpha$ and $\beta$ are inference hyperparameters. This allows DHO to emulate the effects of tuning hyperparameters of SHO without retraining.
OOD generalization.
Although this paper focuses on in-distribution generalization, we also evaluate DHO on ImageNet-based out-of-distribution (OOD) benchmarks [73, 93, 37, 36]. Due to page limits, we present the details in Section D.2; in summary, combining DHO with VPT [43], CoOP [120], and PromptSRC [45] consistently improves their OOD performance.
Training cost and inference efficiency with ToMe.
Table 8 reports the training cost of DHO, and Table 9 shows ToMe [5] significantly reduces inference overhead with minimal performance drop.
@@ 5 Conclusion, Limitation, and Future Work
We presented DHO, a framework that leverages both labeled samples and teacher predictions for knowledge distillation from pretrained VLMs in semi-supervised settings. Through [F1–4], we empirically demonstrated that DHO is a simple yet effective approach. However, our work is currently limited to image classification tasks. Future work could extend DHO to other computer vision and multimodal tasks, such as object detection, segmentation, and language modeling. With suitable architectural adaptations, we believe the dual-head design of DHO can facilitate efficient knowledge transfer from foundation models across a broader range of applications.
@@
  • Theoretical Analysis (Appendix A): provides mathematical foundations and theoretical guarantees for our approach.
@@
  • Algorithms and Implementation (Appendix B): presents detailed pseudocode (Section B.1), implementation specifics (Section B.2), and computational overhead analysis (Section B.3).
@@
  • Datasets (Appendix C): describes the datasets used in our experiments, including statistics and preprocessing details.
@@
  • Additional Experiments (Appendix D): presents MobileNet experiments (Section D.1) and out-of-distribution generalization results (Section D.2).
@@
  • Additional Analyses (Appendix E): contains gradient analysis (Section E.1), non-linear head design studies (Section E.2), and further dual-head investigations (Section E.3).
Appendix A Theoretical Analysis
@@
In this section, we provide a theoretical analysis of our Dual-Head Optimization (DHO) framework. We establish that DHO effectively addresses single-head logit distillation [38, 13] by decoupling conflicting gradients through specialized heads during training. We prove that post-training, the optimal prediction from our dual-head model—formulated as a weighted combination of the heads’ outputs—is mathematically equivalent to the optimal solution of conventional single-head distillation. This equivalence provides theoretical justification for our approach while eliminating gradient conflicts. Furthermore, DHO enables efficient adaptation to various datasets through tunable hyperparameters ($\alpha$ and $\beta$) without requiring model retraining. Note that in this section we slightly abuse the notation of the main paper for clarity, e.g., we denote $p_{\tau}$ as teacher predictions with temperature scaling $\tau$.
A.1 Single-Head Optimization

### v2

@@ 3.3 Analysis
Minimal computational overhead of DHO.
As shown in Tab.˜4, DHO introduces negligible computational overhead with minimal parameter increase on the ImageNet with 1000 classes, which can be further reduced down for datasets with fewer classes ($C<1000$). We further provide inference computational overhead for other models in Tab.˜9.
Figure 6: Results of ablation studies on dual-heads interpolation strategy in Eq.˜8 of DHO.
Figure 7: Qualitative results on challenging cases.
Effectiveness of dual-head interpolation.
We evaluate the effectiveness of dual-head interpolation (Eq.˜8) by comparing DHO with CE+KD (logit), CE+KD (feature), and ablations DHO ($h_{\text{CE}}$) and DHO ($h_{\text{KD}}$), which predict using only one head (i.e., $\alpha=1$ or 0). As shown in Fig.˜6, DHO outperforms DHO ($h_{\text{CE}}$) by an average of 1.6% across 11 datasets, with a maximum gain of +3.4% on ImageNet and no degradation on any dataset. Since $\alpha$ and $\beta$ are inference-time hyperparameters, dual-head interpolation introduces minimal overhead while consistently improving or maintaining performance. We also investigate the effectiveness of our adaptive weighting strategy in Section˜E.4. Fig.˜7 illustrates three challenging examples: CE head ($h_{\text{CE}}$) is correct in the first, KD head ($h_{\text{KD}}$) in the second, and both fail in the third case, yet the proposed combined prediction is correct—demonstrating the ability of DHO to resolve individual head failures. See Section˜F.2 for additional analysis on theses challenging examples.
@@
CE+KD (feature) (Yang et al., 2024a) 62.3 85.0
CE+KD (logit) (Chen et al., 2020) 66.2 88.8
@@
Enhanced feature representation of DHO.
To validate our claim that mitigating gradient conflicts improves feature representations, we evaluate features using the standard linear evaluation protocol (Chen et al., 2020). We train CE+KD (feature), CE+KD (logit), and DHO under the 16-shot semi-supervised setting on ImageNet, freeze the feature extractor $g$, and train a new prediction linear head $h_{\text{LE}}$ on the top of $g$ using fully labeled data. As shown in Section˜3.3, DHO achieves higher Top-1 and Top-5 accuracy than other methods (e.g., +0.9% and 0.5%, respectively). To further assess feature quality, we visualize embeddings $z$ using t-SNE (Van der Maaten & Hinton, 2008) in Fig.˜8. Compared to the CE+KD (logit) baseline, DHO produces more compact and class-separated feature clusters. These results support our claim that DHO enhances feature representations by mitigating gradient conflicts; this improvement leads to better performance of DHO compared to SHO, as discussed through Tab.˜1, Figs.˜4, 5, 11 and 11.
Table 6: Results on language-aware initialization and KD-head alignment for VLM students on ImageNet with 1% labeled data.
Init ($h_{\text{CE}}$$h_{\text{KD}}$). Align. Accuracy (%)
@@
Effectiveness of Init. and Align in Section˜2.3.
To assess the effect of the proposed language-aware initialization and KD-head alignment in Section˜2.3, we run ablations that apply language-aware initialization (Init.) to either $h_{\text{CE}}$ or $h_{\text{KD}}$, and optionally enable KD-head alignment (Align.). We evaluate on ImageNet with 1% labels using a ViT-B/16 student and ViT-L/14 teacher for computational efficiency. As shown in Tab.˜6, applying Init. to either head independently improves accuracy (e.g., +0.2/+0.3%), and adding Align. yields additional gains (e.g., +0.4%).
Figure 9: Grid search results for $\alpha$ and $\beta$.
@@
In Theorem˜1, we show that under mild assumptions (Assumption˜1), DHO $\epsilon$-approximates the SHO baseline, namely CE+KD (logit), in the $\ell_{1}$ norm by setting $\alpha{=}\lambda$ and $\beta{=}1$. This enables DHO to emulate SHO hyperparameter tuning at inference time without retraining. To study how $\alpha$ and $\beta$ affect performance, we visualize a grid search on ImageNet with ResNet-50 under the 16-shot setting. We use ImageNet because it lacks a validation set; thus $\alpha$ and $\beta$ are heuristically set to 0.4 and 0.5 in Section˜3.1. As shown in Fig.˜9, with balanced heads ($\alpha\approx 0.5)$, performance remains stable for $\beta\in[0.1,1]$, and it degrades at extreme $\alpha$ values ($\alpha\approx 0$ or $1$) regardless of $\beta$. The accuracy peaks at 65.5% with balanced heads ($\alpha\approx 0.5$) and a modest temperature ($\beta\approx 0.3$), close to the performance of our heuristic setting (64.7%). Importantly, these gains require no additional training, demonstrating the efficiency of post-training hyperparameter search for the proposed dual-head interpolation in Section˜2.3.
@@ 4 Conclusion, Limitation, and Future Work
We identify the fundamental challenge of transferring the zero-/few-shot capabilities of large vision-language models (VLMs) to task-specific models using limited labeled data in semi-supervised settings. Conventional knowledge distillation (KD) methods suffer from gradient conflicts where teacher signals and labeled data signals interfere with each other. We propose DHO (Dual-Head Optimization), a plug-and-play framework using two classification heads with separate objectives. It mitigates gradient conflicts in both heads and feature extractor, improving feature representations, while enabling flexible post-training hyperparameter adjustment via linear combination of outputs. Experiments across 15 datasets show DHO outperforms conventional KD methods, achieving state-of-the-art results on both ImageNet semi-supervised learning with fewer parameters and on out-of-distribution tasks when combined with existing adaptation techniques.
Limitations and future work.
While our main focus is to address the core problem of gradient conflicts arising from general knowledge of VLMs and task-specific patterns from labeled data, we acknowledge several limitations that present opportunities for future research.
We primarily focus on VLMs as general knowledge source, as they provide strong zero-shot and few-shot capabilities in visual recognition tasks. However, we believe the fundamental problem of gradient conflicts between general foundational knowledge and task-specific objectives extends beyond VLMs. This conflict likely emerges in various scenarios where large pre-trained foundation models (such as instruction-tuned language models (Wei et al., 2021; Liu et al., 2023a; Bai et al., 2023; Achiam et al., 2023; Team et al., 2023)) are adapted to specialized downstream tasks. Exploring how DHO performs across diverse foundation models and modalities remains an important direction for future work.
Our implementation is limited to visual recognition tasks, as they represent the most fundamental domain in computer vision and provide an ideal testbed for analyzing gradient conflicts in knowledge transfer. Also, VLMs’ strong zero-shot and few-shot capabilities in visual recognition tasks make them natural candidates for knowledge distillation. However, extending our approach to more complex visual understanding tasks such as object detection and segmentation would be a promising direction with dedicated architectural adaptations.
Reproducibility Statement
We ensure reproducibility by conducting all experiments on publicly available datasets including ImageNet (Russakovsky et al., 2015), Caltech101 (Fei-Fei et al., 2004), and nine other standard benchmarks detailed in Appendix˜D. All experimental configurations are fully specified in Section˜3.1, including exact hyperparameters, learning rate and optimizer settings, model architectures (ResNet-18, ResNet-50, MobileNetV2, ViT-B/16, ViT-L/14), and training procedures with dual-head optimization detailed in Algorithm 1. We use publicly available pre-trained models (CLIP ResNet-50, CLIP ViT variants from OpenAI, DINO ResNet-50) with exact checkpoint specifications provided in Tab.˜7. The inference procedure with hyperparameters $\alpha$ and $\beta$ is fully documented in Algorithm 2, with specific values for each setting ($\alpha=0.4$ for zero-shot, $\alpha=0.2$ for few-shot teachers, $\beta=0.5$ when validation unavailable). We commit to releasing our complete codebase, training scripts, pretrained checkpoints for ImageNet, and evaluation protocols upon acceptance. All experiments are conducted using PyTorch with fixed random seeds on NVIDIA RTX 4090 GPUs (4× for ImageNet, 8× for VLM distillation, single GPU for other benchmarks).
Ethics Statement
Our work presents no new ethical concerns as DHO is a purely technical contribution for knowledge distillation using existing publicly available datasets (ImageNet, Caltech101, and standard computer vision benchmarks) that contain no personally identifiable information. No additional data collection, human subjects research, or sensitive information processing is involved in this work. We acknowledge that vision-language models may contain biases from their pre-training data, which our distillation framework preserves without amplification. The computational requirements vary by dataset scale (single GPU for most benchmarks, 4× GPUs for ImageNet, 8× GPUs for VLM distillation), which remains modest compared to training large vision-language models from scratch, promoting research accessibility while minimizing environmental impact.
@@
  • Related Work (Appendix˜A) discusses previous work relevant to ours, such as vision-language pre-training, data-limited adaptation of VLMs, knowledge Distillation (KD), and dual head approaches.
@@
  • Theoretical Analysis (Appendix˜B): provides mathematical foundations and theoretical guarantees for our approach.
@@
  • Algorithms and Implementation (Appendix˜C): presents detailed pseudocode (Section˜C.1), implementation specifics (Section˜C.2), and computational overhead analysis (Section˜C.3).
@@
  • Datasets (Appendix˜D): describes the datasets used in our experiments, including statistics and preprocessing details.
@@
  • Additional Experiments (Appendix˜E): presents MobileNet experiments (Section˜E.1), additional results with KD methods (Section˜E.3), additional results with gradient surgery methods (Section˜E.2), additional results with adaptive weighting (Section˜E.4), and results of out-of-distribution generalization with fully-trained models (Section˜E.5).
@@
  • Additional Analyses (Appendix˜F): contains non-linear head design studies (Section˜F.1), and further dual-head investigations (Section˜F.2).
Appendix A Related Work
Vision-language pre-training.
The emergence of vision-language pre-training has marked a significant breakthrough, enabling the use of extensive image-text pairs collected from the web (Wang et al., 2023; Chen et al., 2023) to train powerful vision encoders transferable to various vision tasks (Gan et al., 2022; Zhang et al., 2024b). Early works such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) leveraged contrastive learning techniques to align images and text into a joint representation space, facilitating zero-shot transfer via language prompts. Building on these foundations, subsequent research has focused on improving vision-language models through enhanced training methodologies (Dong et al., 2023; Gao et al., 2022; Yu et al., 2022; Zhai et al., 2023), as well as scaling models and datasets (Yu et al., 2022; Li et al., 2023; Dehghani et al., 2023; Sun et al., 2023a; Cherti et al., 2023; Fang et al., 2023; Sun et al., 2024; Guo et al., 2024) with their zero-shot transfer capabilities (Jia et al., 2021; Zhai et al., 2022; Pham et al., 2023; Liu et al., 2023b). In contrast, our work focuses specifically on target tasks with compact models, aiming to distill knowledge from these large VLMs effectively.
Data-limited adaptation of VLMs.
To preserve pretrained semantic features of VLMs during adaptation with limited data, several approaches have been proposed. Prompt tuning (Lester et al., 2021), initially designed for language models, has been successfully extended to vision tasks. Various methods (Jia et al., 2022; Zhou et al., 2022b; a; Khattak et al., 2023a; Zhu et al., 2023; Khattak et al., 2023b; Menghini et al., 2023; Zhao et al., 2024; Roy & Etemad, 2023; Zhang et al., 2024a; Lafon et al., 2025) have demonstrated the effectiveness of training learnable prompts while keeping the base model frozen. Adapters (Gao et al., 2024; Zhang et al., 2021; Yu et al., 2023b; Silva-Rodriguez et al., 2024) provide an alternative approach by introducing lightweight, trainable modules while maintaining the pre-trained backbone intact. LP++ (Huang et al., 2024) has shown that simple linear layers can effectively adapt CLIP representations in data-limited settings. Note that our work is orthogonal to these approaches: we aim to distill the knowledge of pretrained VLMs into compact models under data-scarce scenarios, making these adaptation methods complementary and applicable to both teacher VLMs in our framework and student models when they are also VLMs.
Knowledge Distillation (KD; Hinton, 2015) enables transferring knowledge from large teacher models to compact student architectures, particularly in data-constrained settings. Researchers have explored synthetic data generation (Lopes et al., 2017; Kimura et al., 2018; Nayak et al., 2019; Yoo et al., 2019; Chen et al., 2019; Yin et al., 2020; Fang et al., 2021a; Nguyen et al., 2022; Patel et al., 2023; Yu et al., 2023a; Liu et al., 2024; Tran et al., 2024; Wang et al., 2024), semi-supervised (Chen et al., 2020; He et al., 2021; Du et al., 2023; Yang et al., 2024b), and unsupervised KD using self-supervised teachers (Fang et al., 2021c; Abbasi Koohpayegani et al., 2020; Navaneet et al., 2021; Wang et al., 2022; Xu et al., 2021; Singh & Wang, 2025). In the VLM domain, recent works (Fang et al., 2021b; Wu et al., 2023; Sun et al., 2023b; Yang et al., 2024a; Vasu et al., 2024; Udandarao et al., 2024; Yang et al., 2024c) distill from large-scale vision-language models into smaller architectures, often using transductive (Kim et al., 2024; Chen et al., 2024) or multi-stage unsupervised strategies (Vemulapalli et al., 2024; Wu et al., 2025; Mistretta et al., 2025). Meanwhile, KD remains challenging due to numerous issues, including model capacity gaps (Cho & Hariharan, 2019; Mirzadeh et al., 2020; Zhu & Wang, 2021; Huang et al., 2022; Li et al., 2024) and inconsistencies between soft and hard targets (Zhang et al., 2023). These challenges are further complicated by misalignment between labeled data and foundational knowledge, especially in few-shot learning scenarios where limited labeled examples may not fully capture the rich semantic understanding of foundation models.
Dual head approaches.
We also consider existing dual-head KD approaches: SSKD (He et al., 2021) uses dual-heads, with each head trained on labeled and unlabeled sets respectively, assuming different data distributions between them, and DHKD (Yang et al., 2024d) introduces binary KD loss working on logits before softmax to prevent neural collapse (Papyan et al., 2020), but both methods inference using only single prediction head $h_{\text{CE}}$. While these previous KD methods adopt dual-head architectures, they do not target distillation from foundation models or combine predictions at inference. In contrast, DHO addresses gradient conflicts in this setting and leverages dual-head aggregation at inference to enhance performance with minimal hyperparameter tuning costs.
Appendix B Theoretical Analysis
@@
In this section, we provide a theoretical analysis of our Dual-Head Optimization (DHO) framework. We establish that DHO effectively addresses single-head logit distillation Hinton (2015); Chen et al. (2020) by decoupling conflicting gradients through specialized heads during training. We prove that post-training, the optimal prediction from our dual-head model—formulated as a weighted combination of the heads’ outputs—is mathematically equivalent to the optimal solution of conventional single-head distillation. This equivalence provides theoretical justification for our approach while eliminating gradient conflicts. Furthermore, DHO enables efficient adaptation to various datasets through tunable hyperparameters ($\alpha$ and $\beta$) without requiring model retraining. Note that in this section we slightly abuse the notation of the main paper for clarity, e.g., we denote $p_{\tau}$ as teacher predictions with temperature scaling $\tau$.
B.1 Single-Head Optimization